An acrobat swings from a trapeze with a frequency of 0.5Hz what is the length of the trapeze cable ?

1
Expert's answer
2022-01-24T11:13:54-0500

We know that

Time period

"T=2\pi\sqrt{\frac{L}{g}}"

"T=\frac{1}{f}"

"\frac{1}{f}=2\pi\sqrt{\frac{L}{g}}"

"L=\frac{g}{4\pi^2 f^2}"

Put value

"L=\frac{9.8}{4\times3.14^2\times 0.5^2}"

L=0.9939m